# Distinguishing Tidal Disruption Events and Changing-look Active Galactic Nuclei via Variation of Mid-infrared Color

### v2

@@ 4.2.2 Influence by dust shape
The pc-scale dust structures in active galaxies generally have torus shape, as indicated by the AGN unified model (e.g. Antonucci, 1993). However, there is no clear evidence for the anisotropy of nuclear dust structure in inactive galaxies. The ANT samples are selected in the optical band, and hence, their toruses are more likely to be face-on. Thus, the difference in the dust geometry can explain why ANTs have smaller ICs than TDEs.
@@ 4.3 Promising explanations
We preferred to explain the smaller ICs in ANTs than in TDEs by differences in dust structure or dust property, because the two types of outbursts have similar UV LCs. Possible explanations include differences in dust chemical composition, in radial optical depth of the dust structure, and in dust geometry. Among them, the difference in dust geometry (face-on torus in ANTs versus spherical shell in TDEs) is supported by observations and hence is promising. If this factor is indeed dominate, we predicted that the IC difference between TDEs and ANTs in MIR samples would be smaller or even disappear, because dusty toruses in MIR selected ANTs should have larger inclinations than optically selected ANTs.
@@ 5 Applying to MIR outbursts
As we have developed a new method to distinguish TDEs/ANTs from CLAGNs, we applied it to samples of MIR outbursts. The samples included the MIRONG sample of Jiang et al. (2021) and the WTP sample of Masterson et al. (2024).
@@ 5.1 MIRONG sample
[…]
The MIR properties of the 106 objects and the 21 objects removed when selecting the rising phase are displayed in Figure 8. The MIR luminosity $L_{\rm max,W2}$ of the sample is in the range of $\sim 10^{42}-10^{44}$ erg s-1, and the rise time $t_{\rm rise}$ is in the range of $<200$ and $>2000$ days. These ranges are similar to those of the optically selected sample, and hence, this sample is suitable for the method distinguishing TDE/ANT and CLAGN developed in section 3.
@@ 5.2 WTP sample
[…]
The MIR properties of the 15 objects and the three excluded objects are displayed in Figure 10. The MIR luminosity $L_{\rm max,W2}$ of the sample is in the range of $\sim 10^{42}-10^{43.5}$ erg s-1, lower than the optically selected sample. This may be because M24 limited the distance to $<$200 Mpc for the sample. The rise time $t_{\rm rise}$ is in the range of $<200$ and $\sim 1000$ days, shorter than the optically selected sample. This is because M24 required a rapid rise in the MIR LC.
@@ 5.3 Promising TDE candidates
SDSS J0842+2357 0.06353 MIRONG(SF) -- $0.74\pm 0.11$ 12.80 0.81 42.60 $<154$
SDSS J1549+3327 0.08565 MIRONG(SF) -- $0.75\pm 0.03$ 12.68 2.21 43.17 $<160$
WTP14adbjsh 0.0106 WTP(Gold) -- $0.74\pm 0.03$ 7.99 2.10 43.17 $<182$
We selected MIR outbursts in MIRONG and WTP samples that have a probability of falling into U region $p_{U}$ of $>70\%$, and are hosted in non-Seyfert galaxies. There are eight outbursts, seven from the MIRONG sample and one from the WTP sample. We list the MIR information of these TDE candidates in Table 3.
@@ 6.2 Fractions of TDEs, ANTs and CLAGNs in MIR selected samples
Can CLAGNs occur in non-Seyfert galaxies? Non-Seyfert galaxies were selected using narrow-line-ratio and MIR color diagnoses (Jiang et al., 2021; Masterson et al., 2024). These diagnoses cannot fully ruled out weak AGNs whose signals are overwhelmed by star formation. Thus, for the outbursts in the MIRONG and WTP non-Seyfert samples that were diagnosed as CLAGNs by MIR color, their host galaxies may have weak AGNs, even though there was no clear AGN signal before the outbursts.
The fraction of TDEs in the WTP sample (62%–82%) is higher than that in the MIRONG non-Seyfert sample (45%–63%). This is possibly because Masterson et al. (2024) required an additional criterion that the MIR LC must have a fast rise and a slow, monotonic decay. However, CLAGNs could not be completely excluded after adding this criterion, as the fraction of CLAGNs in the WTP sample is still $\gtrsim 20\%$, possibly because some CLAGNs also meet this criterion. In addition, it is likely that some TDEs, for example, AT 2019qiz (Figure A2) whose MIR emission rises nonmonotonically in $>925$ days, fail to pass this criterion. Therefore, this standard may improve the purity of TDEs while losing completeness.
@@ 6.3 Implications and future works
Jiang et al. (2021) and Masterson et al. (2024) obtained incident rates of MIR outbursts with their samples of $5.4\times 10^{-5}$ and $2.0\pm 0.3\times 10^{-5}$ galaxy-1 yr-1, respectively. We estimated the incident of IR TDEs using their results and our estimations of the TDE fractions. For the MIRONG sample, with the 77 MIR outbursts in non-Seyfert galaxies that are not supernovae and are radio-quiet, we could estimate a rate of $3.4\times 10^{-5}$ galaxy-1 yr-1 considering that non-Seyfert galaxies make up $\sim 90\%$ of SDSS galaxies (Kewley et al., 2006). Assuming that 45%–63% of them are TDEs, the incident rate of IR TDEs is $1.5-2.1\times 10^{-5}$ galaxy-1 yr-1. For the WTP sample, Masterson et al. (2024) estimated the rate using 12 MIR outbursts in their gold sample. According to our estimation, 62%–82% of their total 18 outbursts are TDEs, and hence the rate should be modified to $1.6-2.8\times 10^{-5}$ galaxy-1 yr-1, which is consistent with the rate Masterson et al. (2024) derived using only their gold sample. Therefore, the results from the MIRONG and WTP samples are consistent, and the final rate is $1.5-2.8\times 10^{-5}$ galaxy-1 yr-1, comparable with the optical TDE rate of $3.2^{+0.8}_{-0.6}\times 10^{-5}$ galaxy-1 yr-1 (Yao et al., 2023) within the margin of error.
@@
Part of IR TDEs are bright in both optical and IR bands and are also included in optical TDEs. Jiang et al. (2021) and Masterson et al. (2024) estimated optical detections for MIR outbursts of 11% and 22%, respectively. In addition, of the 33 optical TDEs in Yao et al. (2023), only two (AT 2019dsg and AT 2019qiz) are MIR bright. Thus, there is not much overlap between optical TDEs and IR TDEs. Combining results from optical, X-ray and IR surveys, we estimated a total TDE rate of $\sim 5-8\times 10^{-5}$ galaxy-1 yr-1.
@@
So far, our method has not been able to classify most of the MIR outbursts clearly, mainly due to the large measurement errors of MIR colors and the semi-annual sampling interval of the WISE project, which is not frequent enough. Nevertheless, some MIR outbursts can already be classified based on the available data. For example, SDSS J154843.06+220812.6, whose radio counterpart VT J154843.06+220812.6 has been studied in detail by Somalwar et al. (2022), is a likely CLAGN because of its CVR of $-0.08\pm 0.03$, corresponding a $p_{L}$ of 100%.
@@
Although the WISE telescope was retired in July 2024, the Near-Earth Object Surveyor (Mainzer et al., 2023) will launch in 2027. Furthermore, there will be some wide-field NIR survey projects in the future, such as the wide-field infrared transient explorer (WINTER, Lourie et al., 2020), PRime-focus Infrared Microlensing Experiment (PRIME, Kondo et al., 2023), and the Nancy Grace Roman Space Telescope (Akeson et al., 2019). These projects will discover a large amount of IR outbursts in galaxy centers, and obtain their high-quality, high-cadence NIR LCs. Our simulations predict that the NIR color $H-K$ of TDEs also turns red faster in the rising phase than CLAGNs, with a pattern similar to MIR color (Appendix E.3). If true, IR outbursts can also be classified based on the NIR CVR. Due to the high quality and cadence of NIR data, we expected to obtain more IR outbursts with precise classification by the NIR CVR.
@@ 7 Summary and conclusions
Using simulations involving the dust radiative transfer model, we explored the origins of the two phenomena. The MIR color of TDEs and ANTs turning red quickly possibly results from their UV LCs characterized by “rapid rise, short peak and long tail”. In addition, the difference in IC between TDEs and ANTs may be related to the difference in the geometry of the echoing dust: the dust structure around TDEs is approximately spherically symmetric, while that around ANTs has a torus shape with face-on inclination.
@@
Using this method, we classified 127 MIR outbursts in the MIRONG sample of Jiang et al. (2021) and 18 in the WTP sample of Masterson et al. (2024). For the MIRONG sample, we divided it into a non-Seyfert subsample and a Seyfert subsample, according to the types of host galaxies, containing 77 and 50 MIR outbursts, respectively. We found that in all three samples, both likely TDEs/ANTs and likely CLAGNs exist, and hence, they are all mixtures of TDEs/ANTs and CLAGNs. We selected 11 promising TDE candidates based on a high probability of CVR falling in the U region or a short rise time. We examined our classifications with those in the literature based on follow-up optical spectra and X-ray observations and found no apparent contradiction.
@@
This work is supported by the National Natural Science Foundation of China (NFSC, 12103002) and the University Annual Scientific Research Plan of Anhui Province (2022AH010013, 2023AH030052). This publication makes use of data products from NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the Planetary Science Division of the National Aeronautics and Space Administration.
@@ Appendix B K-correction for the MIR color
Figure B1: (a): Rest-frame $W1-W2$ color as functions of temperature for different types of SEDs. Blue, red and green lines represent results of optically thin graphite and silicate dust, and blackbody. For graphite and silicate dust, results assuming grain radii of 0.0316, 0.1 and 0.316 $\mu$m are shown using solid, dotted and dashed lines. (b): The K-correction for $z=0.1$ as functions of color in the observer’s frame for different types of SEDs. The meanings of colors and styles of the lines are the same as panel (a). The grey shades show the allowed range of K-correction, and the black solid line shows the final K-correction curve. (c): The final K-correction curves and allowed range for $z=0.1$, 0.2 and 0.3.
The color $C=W1-W2$ of a target in the observer’s frame is determined by SED and redshift. We considered three types of SEDs: blackbody, optically thin graphite, and silicate. The blackbody SED is expressed as:
@@
where $T_{d}$ and $M_{d}$ are the temperature and mass of the dust, and $\kappa_{\lambda}(a)$ is the dust mass absorption coefficient for grain radius $a$. We adopted the absorption coefficients from Laor & Draine (1993) and converted them to $\kappa_{\lambda}$ following Fox et al. (2010).
@@
In Figure B1(a), we show the relationships between rest-frame colors and temperature for different SEDs. The results for grain radius $a=0.0316$, 0.01 and 0.316 $\mu$m are displayed for optically thin graphite and silicate. At the same temperature, graphite dust is the bluest, followed by silicate dust, and blackbody is the reddest. The color of actual dust may be more complex. If the dust is a mixture of graphite and silicate, the color would be between pure graphite and silicate. In addition, if the MIR optical depth of the dust is on the order of magnitude of 1, the color would be between those of optically thin dust and blackbody. Therefore, despite the complexity of the SED of actual dust, its color can still be constrained by the above simple SED models.
For different redshifts, temperatures, and types of SEDs, we calculated K-correction $K(z)=C(z)-C(0)$, where $C(z)$ and $C(0)$ are colors in the observer’s frame and the rest frame, respectively. We show the $K(z)$ as a function of $C(z)$ for $z=0.1$ in Figure B1(b). We found that the functions assuming different types of SEDs are approximately a series of parallel lines. For $z<0.35$, the top is always graphite with $a=0.316$ $\mu$m, and the bottom is always the blackbody. Therefore, we adopted the average of the $K(z)$ given by these two types of SEDs as the final K-correction curve (the black line), and estimated the systematic error by half of their difference. The final K-corrections and the systematic errors both increase as $z$ increases, as shown in Figure B1(c). The systematic error caused by uncertain types of SEDs can be roughly estimated as $\sigma_{\rm sys}(K)\approx 0.04z$.
@@ Appendix C Analysis of the 2020 outburst of Mrk 1018
[…]

### v3

@@ 4.2.3 Spherical shell or torus
The ANTs in our sample are selected in the optical band, and hence, their tori are more likely to be face-on. While for TDEs, the dust geometry might be better described by a spherical shell since there is no clear evidence for the anisotropy of nuclear dust structure in inactive galaxies. Therefore, the difference in the dust geometry can explain why ANTs have smaller ICs than TDEs.
In addition, face-on torus results in a slightly larger CVR than the spherical shell with a difference of $\sim 0.1-0.2$ mag yr-1. The CVR difference may be because the echo produced by the peak luminosity is bluer due to smaller emergent optical depth, and has a larger color difference with the echo from outer region with red color.
@@ 4.3 Promising explanations
We preferred to explain the smaller ICs in ANTs than in TDEs by differences in dust property or dust structure, because the two types of outbursts have similar UV LCs. The difference in dust properties requires that the echoing dust of ANTs, i.e., the inner region of an AGN dusty torus, has a higher fraction of graphite and larger grain size than the echoing dust of TDEs. Whereas the difference in dust geometry requires that the echoing dust in ANTs is a face-on torus, while the echoing dust in TDEs is a spherical shell. If the difference in dust geometry is the dominant factor, we predicted that the IC difference between TDEs and ANTs in MIR samples would be smaller or even disappear, because dusty tori in MIR-selected ANTs should have larger inclination angles than optically selected ANTs.
[…]
@@ 5 Applying to MIR outbursts
As we have developed a new method to distinguish TDEs/ANTs from CLAGNs, we applied it to samples of MIR outbursts, including the mid-infrared outbursts in nearby galaxies (MIRONG) sample of Jiang et al. (2021b), and the sample of Masterson et al. (2024), which is part of the Wise Transient Pipeline (WTP) sample.
@@ 5.1 MIRONG sample
The MIRONG sample contains 137 MIR outbursts. We removed four supernovae and four objects with strong radio emissions that may be associated with a jet, all identified by Jiang et al. (2021b). We also removed two objects that contain problematic data points. We reduced the data for the remaining 127 objects and selected quiescent and outburst states and the rising phase as we did for the optically selected sample. The only difference is that we did not require $m_{\rm min,W2}<12.5$ or $\delta{\rm W2}>0.5$ here. In this process, we removed 12 objects whose rising phase began at the first data point of NEOWISE-R. We also removed nine objects with only one data point in the rising phase. Finally, there are 106 objects left for the following analysis.
Figure 10: MIR properties of the MIRONG sample. The upper panel shows $m_{\rm min,W2}$ and $\delta{\rm W2}$, and the lower panel shows $L_{\rm max,W2}$ and $t_{\rm rise}$. Star-forming, Composite, LINER and Seyfert galaxies are shown in red, green, orange and blue symbols. Note that the objects in Jiang et al. (2021b) with broad emission lines (Seyfert 1) and those with Seyfert-like narrow line ratios (Seyfert 2) were both treated as Seyfert galaxies in this work. The meanings of symbols (dot, cross and inverted triangle) are the same as Figure 3.
The MIR properties of the 106 objects and the 21 objects removed when selecting the rising phase are displayed in Figure 10. The MIR luminosity $L_{\rm max,W2}$ of the sample is in the range of $\sim 10^{42}-10^{44}$ erg s-1, and the rise time $t_{\rm rise}$ is in the range of <200<200 and >2000>2000 days. These ranges are similar to those of the optically selected sample, and hence, this sample is suitable for the method distinguishing TDE/ANT and CLAGN developed in section 3.
@@ 5.2 WTP sample
Figure 12: MIR properties of the WTP sample. The upper panel shows $m_{\rm min,W2}$ and $\delta{\rm W2}$, and the lower panel shows $L_{\rm max,W2}$ and $t_{\rm rise}$. The gold and silver subsamples are shown in red and blue symbols. The meanings of symbols (dot, cross and inverted triangle) are the same as Figure 3.
The MIR properties of the 15 objects and the three excluded objects are displayed in Figure 12. The MIR luminosity $L_{\rm max,W2}$ of the sample is in the range of $\sim 10^{42}-10^{43.5}$ erg s-1, lower than the optically selected sample. This may be because M24 limited the distance to <<200 Mpc for the sample. The rise time $t_{\rm rise}$ is in the range of <200<200 and $\sim 1000$ days, shorter than the optically selected sample. This is because M24 required a rapid rise in the MIR LC.
@@ 5.3 Promising TDE candidates
SDSS J0842+2357 0.06353 MIRONG(SF) - $0.74\pm 0.11$ - 12.80 0.81 42.60 <154<154
SDSS J1549+3327 0.08565 MIRONG(SF) - $0.75\pm 0.03$ - 12.68 2.21 43.17 <160<160
WTP14adbjsh 0.0106 WTP(Gold) - $0.74\pm 0.03$ - 7.99 2.10 43.17 <182<182
We selected MIR outbursts in MIRONG and WTP samples that have a probability of falling into U region $p_{U}$ of $70\%$, and are hosted in non-Seyfert galaxies. There are eight outbursts, seven from the MIRONG sample and one from the WTP sample. We list the MIR information of these TDE candidates in Table 3.
@@ 6.2 Fractions of TDEs, ANTs and CLAGNs in MIR selected samples
Can CLAGNs occur in non-Seyfert galaxies? Non-Seyfert galaxies were selected using narrow-line-ratio and MIR color diagnoses (Jiang et al., 2021b; Masterson et al., 2024). These diagnoses cannot fully ruled out weak AGNs whose signals are overwhelmed by star formation. Thus, for the outbursts in the MIRONG and WTP non-Seyfert samples that were diagnosed as CLAGNs by MIR color, their host galaxies may have weak AGNs, even though there was no clear AGN signal before the outbursts.
The fraction of TDEs in the WTP sample (62%–82%) is higher than that in the MIRONG non-Seyfert sample (45%–63%). This is possibly because Masterson et al. (2024) required an additional criterion that the MIR LC must have a fast rise and a slow, monotonic decay. However, CLAGNs could not be completely excluded after adding this criterion, as the fraction of CLAGNs in the WTP sample is still $\gtrsim 20\%$, possibly because some CLAGNs also meet this criterion. In addition, it is likely that some TDEs, for example, AT 2019qiz (Figure A2) whose MIR emission rises nonmonotonically in >925>925 days, fail to pass this criterion. Therefore, this standard may improve the purity of TDEs while losing completeness.
@@ 6.3 Implications and future works
Jiang et al. (2021b) and Masterson et al. (2024) obtained incident rates of MIR outbursts with their samples of $5.4\times 10^{-5}$ and $2.0\pm 0.3\times 10^{-5}$ galaxy-1 yr-1, respectively. We estimated the incident of IR TDEs using their results and our estimations of the TDE fractions. For the MIRONG sample, with the 77 MIR outbursts in non-Seyfert galaxies that are not supernovae and are radio-quiet, we could estimate a rate of $3.4\times 10^{-5}$ galaxy-1 yr-1 considering that non-Seyfert galaxies make up $\sim 90\%$ of SDSS galaxies (Kewley et al., 2006). Assuming that 45%–63% of them are TDEs, the incident rate of IR TDEs is $1.5-2.1\times 10^{-5}$ galaxy-1 yr-1. For the WTP sample, Masterson et al. (2024) estimated the rate using 12 MIR outbursts in their gold sample. According to our estimation, 62%–82% of their total 18 outbursts are TDEs, and hence the rate should be modified to $1.6-2.8\times 10^{-5}$ galaxy-1 yr-1, which is consistent with the rate Masterson et al. (2024) derived using only their gold sample. Therefore, the results from the MIRONG and WTP samples are consistent, and the final rate is $1.5-2.8\times 10^{-5}$ galaxy-1 yr-1, comparable with the optical TDE rate of $3.2^{+0.8}_{-0.6}\times 10^{-5}$ galaxy-1 yr-1 (Yao et al., 2023) within the margin of error.
@@
Part of IR TDEs are bright in both optical and IR bands and are also included in optical TDEs. Jiang et al. (2021b) and Masterson et al. (2024) estimated optical detections for MIR outbursts of 11% and 22%, respectively. In addition, of the 33 optical TDEs in Yao et al. (2023), only two (AT 2019dsg and AT 2019qiz) are MIR bright. Thus, there is not much overlap between optical TDEs and IR TDEs. Combining results from optical, X-ray and IR surveys, we estimated a total TDE rate of $\sim 5-8\times 10^{-5}$ galaxy-1 yr-1.
@@
So far, our method has not been able to classify most of the MIR outbursts clearly, mainly due to the large measurement errors of MIR colors and the semi-annual sampling interval of the WISE project, which is not frequent enough. Nevertheless, some MIR outbursts can already be classified based on the available data. For example, SDSS J154843.06+220812.6, whose radio counterpart VT J154843.06+220812.6 has been studied in detail by Somalwar et al. (2022), is a likely CLAGN because of its CVR of $0.08\pm 0.03$, corresponding a $p_{L}$ of 100%.
@@
Although the WISE telescope was retired in July 2024, the Near-Earth Object Surveyor (Mainzer et al., 2023) will launch in 2027. Furthermore, there will be some wide-field NIR survey projects in the future, such as the wide-field infrared transient explorer (WINTER, Lourie et al., 2020), PRime-focus Infrared Microlensing Experiment (PRIME, Kondo et al., 2023), and the Nancy Grace Roman Space Telescope (Akeson et al., 2019). These projects will discover a large amount of IR outbursts in galaxy centers, and obtain their high-quality, high-cadence NIR LCs. Our simulations predict that the NIR color HKH-K of TDEs also turns red faster in the rising phase than CLAGNs, with a pattern similar to MIR color (Appendix E.2). If true, IR outbursts can also be classified based on the NIR CVR. Due to the high quality and cadence of NIR data, we expected to obtain more IR outbursts with precise classification by the NIR CVR.
@@ 7 Summary and conclusions
Using simulations involving the dust radiative transfer model, we explored the origins of the two phenomena. One phenomenon, the MIR color of TDEs and ANTs turning red quickly, is possibly because of their UV LCs characterized by “rapid rise, short peak and long tail”, and/or because of no or weak contribution from underlying AGN. The other one, ANTs showing bluer initial colors than TDEs, is possibly because of higher fraction of graphite and larger grain size of the echoing dust, and/or because of the difference in dust geometry: dust around ANTs has a torus shape with face-on inclination, while that around TDEs is approximately spherically symmetric.
@@
Using this method, we classified 127 MIR outbursts in the MIRONG sample of Jiang et al. (2021b) and 18 in the WTP sample of Masterson et al. (2024). For the MIRONG sample, we divided it into a non-Seyfert subsample and a Seyfert subsample, according to the types of host galaxies, containing 77 and 50 MIR outbursts, respectively. We found that in all three samples, both likely TDEs/ANTs and likely CLAGNs exist, and hence, they are all mixtures of TDEs/ANTs and CLAGNs. We selected 11 promising TDE candidates based on a high probability of CVR falling in the U region or a short rise time. We examined our classifications with those in the literature based on follow-up optical spectra and X-ray observations and found no apparent contradiction.
@@
This work is supported by the National Natural Science Foundation of China (NFSC, 12103002) and the University Annual Scientific Research Plan of Anhui Province (2022AH010013, 2023AH030052). This publication makes use of data products from NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the Planetary Science Division of the National Aeronautics and Space Administration. This research has made use of the NASA/IPAC Infrared Science Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration, the archival data are DOI: 10.26131/IRSA535 (IRSA, 2022) and DOI: 10.26131/IRSA144 (NEOWISE Team, 2020). Software: This work made use of the following python packages: Numpy (Harris et al., 2020), Matplotlib (Hunter, 2007), Scipy (Virtanen et al., 2020) Astropy (Astropy Collaboration et al., 2022), and emcee (Foreman-Mackey et al., 2013).
@@ Appendix B K-correction for the MIR color
[…]
The color C=W1W2C=W1-W2 of a target in the observer’s frame is determined by SED and redshift. We considered three types of SEDs: blackbody, optically thin graphite, and silicate. The blackbody SED is expressed as:
@@
where $T_{d}$ and $M_{d}$ are the temperature and mass of the dust, and $\kappa_{\lambda}(a)$ is the dust mass absorption coefficient for grain size $a$. We adopted the absorption coefficients from Laor & Draine (1993) and converted them to $\kappa_{\lambda}$ following Fox et al. (2010).
@@
[…]
@@ Appendix C Analysis of the 2020 outburst of Mrk 1018
We refit the same data using the GE model. We reduced the ZTF and ATLAS data and generated host-subtracted LCs following Wang et al. (2024). We fit the data simultaneously instead of fitting ZTF and ATLAS data separately as Lu et al. (2025) did. We fit by the Monte-Carlo Markov-Chain (MCMC) method with the code $\it emcee$} assuming log-uniformly prior distributions for $\sigma$ and $\tau$. The posterior probability density distributions of $\sigma$ and $\tau$ are displayed in Figure C1(a). As can be seen from the figure, the observed data can be reproduced by two different groups of parameters, which can roughly be divided by the threshold of $\sigma=60$ day. For one group $\sigma=33\pm 14$ day and $\tau=116\pm 6$ day, representing fast rising models with shorter rise time than fall time (Figure C1(b)). And for the other group $\sigma=98^{+8}_{-11}$ day and $\tau=83^{+14}_{-10}$ day, representing slowly rising models with similar rise time and fall time (Figure C1(c)). The integral posterior probabilities for the two groups are 17% and 83%, respectively, indicating that slowly rising models have a higher probability, but fast-rising models are also acceptable by the data. Taking into account systematic errors due to AGN intrinsic variability, the differences between the two groups of models would be even smaller. Therefore, we concluded that the LC data alone cannot rule out the possibility of UV LCs with fast rise and slow fall.